Circular No. 8974

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 R4 (KORLEVIĆ) G. V. Williams, Minor Planet Center, has identified recovery obser-

vations of comet P/1999 WJ_{7} (cf. IAUC 7368) on Sept. 3 in incidental astrometry obtained with the Spacewatch 0.9-m telescope by T. H. Bressi. A request by B. G. Marsden to J. V. Scotti for confirming observations yielded images with the same telescope on Sept. 9.4, in which Scotti notes that the comet is slightly diffuse with a 0′.14 tail in p.a. 201^{o}. The indicated correction to the prediction on MPC 59598 is ΔT = −0.3 day.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.3.24478  0^{h}09^{m}59.^{s}68  +4^{o}59′39′′.4  20.3

The new astrometry, together with the following revised orbital elements and an ephemeris, appear on MPEC 2008-R53.

Epoch = 2010 Feb. 13.0 TT

T = 2010 Feb. 8.1984 TT  ω = 154.^{o}5443
e = 0.315128  Ω = 290.5650 }2000.0
q = 3.182125 AU  i = 2.9758

a = 4.646302 AU n^{o} = 0.0984108 P = 10.02 years

COMET P/2008 R5 (LINEAR-NEAT) Williams has also identified recovery observations of comet P/2001

TU_{80} (cf. IAUC 7753) on Sept. 8 in incidental astrometry obtained with the Spacewatch 1.8-m telescope by Scotti. Again, a request to Scotti for confirmation yielded images with the 0.9-m reflector on Sept. 9.4 that show a coma diameter of 6′′ and a faint tail extending 0′.16 in p.a. 225^{o}. The indicated correction to the prediction on MPC 54171 is ΔT = −0.3 day.

2008 UT  α_{2000}  δ_{2000}
Sept.8.43387  6^{h}13^{m}37.^{s}98  +18^{o}59′47′′.5

The new astrometry, together with the following revised orbital elements and an ephemeris, appear on MPEC 2008-R54.

Epoch = 2008 Nov. 30.0 TT

T = 2008 Dec. 9.2712 TT  ω = 355.^{o}0388
e = 0.470561  Ω = 109.1065 }2000.0
q = 1.940235 AU  i = 6.5812

a = 3.664697 AU n^{o} = 0.1404905 P = 7.02 years

2008 September 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green